Determine the extrema of gew v= bu- Subject to the constraint Kitys 360
Answer #1

2 According to questron, function goven gruan) = bu .bt subject to constant 22tog? = 360 Both the functron and the constant aلأولى n² + y2 = 360 n of rom equation all 2 + *2=360 5 CA ) H, 20 we put in for yand above equatron 72 =360 - Ea2=360 4 = n²